A student dropped few pieces of marble in dilute hydrochloric acid, contained in a test tube. The evolved gas was then passed through lime water.what change would be observed in lime water. What will happen if excess of gas is passed through lime water. Write balanced chemical equations for all the changes observed.
1
Expert's answer
2017-05-29T09:46:10-0400
Marble consists of calcium carbonate CaCO3.

CaCO3 + 2HCl = CaCl2 + CO2 + H2O

The gas is CO2.

Lime water is a saturated water solution of Ca(OH)2.

Ca(OH)2 + CO2 = CaCO3 + H2O

The lime water will become cloudy.

In case of excess of the gas, whole amount of calcium hydroxide (Ca(OH)2) will turn to CaCO3.